Calculate the quantity of electricity pass when a current of 0.45 amps Flows for 25mins
1
Expert's answer
2020-06-11T10:29:20-0400

The quantity of electricity is measured in coulombs C. It is calculated using the formula below.


"Q=I*t" where,

Q is the quantity of electricity in C,

I is the current in amperes (amps) and

t is time in seconds.


In our question we are given,

I=0.45 amps and

t=25 minutes


Therefore we use the formula "Q=I*t" to find Q

First convert 25 min into seconds


"t=25*60=1500" seconds


"Q=0.45*1500"

"=675C"